OpenStudy (anonymous):

How do you use the formula square root 24d ?

OpenStudy (ineedserioushelp):

by 24d do you mean distance?

OpenStudy (ineedserioushelp):

like in the formula s = 24d, for speed equals 24 time distance?

OpenStudy (anonymous):

yes the question where the formula was Using the formula, square root 24d, find the speed of a car in miles per hour if it took 150 feet to stop. Round your answer to a whole number.

OpenStudy (ineedserioushelp):

Well let's say a bicycle is going to the store, and it took him 150 feet until he got there, how fast was he going? s = sqrt (24d) s^2 = 24 * 150 s^2 = 3600 s = 60 So, we used the formula to find out how fast he was going.

OpenStudy (anonymous):

Yes. Thank you. I also wanted to know how to use the formula? How do I order the numbers in the expression and what do I do?

OpenStudy (ineedserioushelp):

Well the basic formula is: s = square root (24d) substitute d for the distance. so let's say the distance is 150 like in your problem we will use s^2 instead of s = square root so, s^2 = 24 * 150 s^2 = 3600 √3600 = 60 so s = 60

OpenStudy (anonymous):

ok thanks so much. Our teacher hadn't even covered this yet. Thanks again.

OpenStudy (ineedserioushelp):

No problem :)

OpenStudy (ineedserioushelp):

If you could leave me a medal I'd appreciate it :))

OpenStudy (anonymous):

I'm new. How do I leave a medal?